# An Exploratory Study on Human-Robot Interaction using Semantics-based Situational Awareness

### v1

@@ Abstract
In this paper, we investigate the impact of high-level semantics (evaluation of the environment) on Human-Robot Teams (HRT) and Human-Robot Interaction (HRI) in the context of mobile robot deployments. Although semantics has been widely researched in AI, how high-level semantics can benefit the HRT paradigm is underexplored, often fuzzy, and intractable. We applied a semantics-based framework that could reveal different indicators of the environment (i.e. how much semantic information exists) in a mock-up disaster response mission. In such missions, semantics are crucial as the HRT should handle complex situations and respond quickly with correct decisions, where humans might have a high workload and stress. Especially when human operators need to shift their attention between robots and other tasks, they will struggle to build Situational Awareness (SA) quickly. The experiment suggests that the presented semantics: 1) alleviate the perceived workload of human operators; 2) increase the operator’s trust in the SA; and 3) help to reduce the reaction time in switching the level of autonomy when needed. Additionally, we find that participants with higher trust in the system are encouraged by high-level semantics to use teleoperation mode more.
@@ I Introduction
The increasing use of robots is engendering challenges and higher demands for Human-robot collaboration. Human-Robot Teams (HRT) need to work together and leverage complementary strengths to achieve the shared mission goal by combining perception capabilities (remote sensing, object detection) and cognition (e.g. scene understanding, decision-making, task planning). Situational Awareness (SA) [1] is one of the significant factors in bridging cognition and perception. Explainable and comprehensive SA can directly benefit robots and humans in many scopes, e.g. task planning [2], variable autonomy [3], and Human-Robot Interaction (HRI).
@@
Semantic information, which is one of the components of SA, is suggested to alleviate the above difficulties [7]. Semantic information is widely discussed in the artificial intelligence community, e.g. in computer vision. However, this literature is predominantly limited to relatively low-level semantics (e.g. image segmentation). The impact of high-level semantics, which requires more understanding and inference, has been less well explored.
@@
Our previous works categorize environmental semantics in the scope of disaster response missions [8] and propose a framework [9] that applies the above categorization and quantifies different understandings of the situation by semantic indicators. Moreover, the framework enables providing robots with an indicator of the overall semantic richness of a scene. This method translates sensor readings, measurements, and detections into a semantics richness score, which can inform remote human operators if the situation requires more attention.
@@
This paper provides new insights into HRT and HRI with respect to high-level semantics. We present three main contributions: i) identifying the usefulness of the proposed semantic richness framework [9]; ii) identifying how high-level semantics affect HRT and HRI in the context of a SAR mission; iii) insights about patterns and relationships among HRT factors and high-level semantics, e.g. Level of Autonomy (LoA) preference versus trust in a variable autonomy [10] system.
@@ II Related work
A wide range of studies explore ways of improving SA from the perspectives of visual design, workload and stress management, human training, and Level of Autonomy (LoA).
Consequently, HRI can be categorized [11, 12] into cognitive responses [1], emotional responses [13], behavioral responses [14], and physiological responses [15, 16]. Our work can partly be considered from a similar perspective, exploring how high-level semantics can enhance SA and benefit HRT and HRI, improving the likelihood of mission success, while reducing human workload.
@@
Figure 2: Test arena for trials without semantic indicators. Note that we need to use two different arenas layout for the two different trials, to minimise the confounder of learning effects between trials.
The above-mentioned studies investigated low-level semantics, but did not explore higher-level semantics. In contrast, our work explores high-level semantics, and its effects on human attention in HRT. [7] evaluated the effects of four interfaces on human stress and workload in a simulated fire surveillance and extinguishing mission. Virtual Reality (VR) interfaces were shown to enhance SA and reduce workload compared to conventional interfaces. The authors used high-level semantics, e.g. displaying information about fire risk, during the experiments. However, they investigated and compared the effects of different interface designs for displaying such data. In contrast, our work seeks to evaluate the impact of the high-level semantic data itself, i.e. performance with versus without displaying high-level semantics.
Other work explores how human training [20] and expertise affect SA. [21] investigated HRT under limited SA conditions, where operators lacked a bird’s-eye view of the environment. Participants were required to navigate robots until they met up together in the experiments. The authors emphasized the importance of training. Expert operators (several hours of training) outperformed others, by minimizing the number of commands issued to the robots. This highlights the critical role of prior knowledge and training in recovering and maintaining SA.
The LoA plays an important role in affecting SA as well. [22] explored human factors in HRT and found that high robot autonomy could reduce human SA. This makes it difficult for humans to regain SA when taking back control of the robot, based on SAGAT questionnaires [23]. The authors recommended a mixed-initiative approach that keeps humans in the loop and balances workload and SA between humans and robots. These related studies help us identify potential confounding factors and minimize their impacts in our experiment design.
@@
III Exploratory study of HRI and HRT with high-level semantics
This paper explores the effects of presenting high-level semantics to a human operator during HRT deployments. Specifically, we explore the performance variance of HRT, with and without the display of semantic indicators (high-level semantics). We also explore the practical utility of our previously proposed framework for real-time computing of such indicators [9].
Our semantics framework [9] quantifies the richness of low-level semantics (salient features that can be understood by a human’s direct intuition) and high-level semantics (evaluation, estimation, or inference of a human’s perspective) [8] in the current scene. These high-level semantic indicators can be used to help a human operator achieve SA. In our experiments, high-level semantics include: signs of human activities; radiation levels, and navigational laser noise. SHA is detected as a distribution of personal belongings. Radiation level is a mapping from radiation sensor readings that enables human operators to visualise radiation intensity. The laser noise refers to the robot’s LiDAR. It helps robots and humans understand whether noise is becoming significant enough to disable the robot’s navigation. Each of these metrics carries a normalized score between 0 and 1, helping assess potential risks or challenges in real time.
In this exploratory study, we use a simulated SAR mission with human participants. Specifically, we simulate the first HRT deployment in a disaster response mission. The HRT needs to explore the environment and find victims while avoiding risks to robots. The robot can operate in two different LoAs: teleoperation and autonomous navigation modes. Humans can switch the LoA dynamically in a Human-Initiative paradigm [24] and can take over teleoperative control of the robot when they feel it is necessary. Moreover, parallel cognitive tasks are assigned to human operators to simulate non-robotic emergencies that may require human attention. We have two major research questions in this experiment: i) how do high-level semantics affect the HRT and HRI, and ii) what is the correlation among HRT factors when affected by high-level semantics?
@@ III-A Arena
The experimental arena is an outdoor post-disaster environment, see Fig. 1 and Fig. 2. The robot is set to autonomously navigate along a pre-planned route, but the human operator can switch to teleoperation mode at any time, and explore other regions away from this route. Two human victims, one radiation source, and one laser noise zone are set up in the arena. Specifically, one human victim can be found by the camera (on the scheduled path, without the robot needing to deviate) by locating the personal belongings. Another victim is hidden behind obstacles, away from the scheduled path. The radiation source and laser noise zones are set to occur at pre-set regions along the planned path. The starting point and endpoint for robots are pre-defined. Note that we use a “within groups” experimental design, where each participant undertakes both experimental conditions (with and without semantic indicators). Therefore, we need to use a different test arena layout for each condition, to minimise learning effects as a confounder. However, the length and complexity of the tasks, and the types of semantics encountered, are kept very similar to ensure meaningful comparability.
@@ III-B Mission and Tasks
We assume that the HRT needs to conduct its first deployment to investigate the site in a SAR mission. The participant can switch back and forth between the two LoAs (teleoperation and autonomy) whenever they choose. In autonomy mode, the robot will try to follow the pre-set route. Three tasks are assigned to the HRT: A) finding victims, B) navigating the robot, and C) parallel task. Task A requires the HRT to find as many victims as they can in the arena. Task B requires the HRT to navigate the robots from a pre-set starting point to the endpoint within set time limits. Task C is a non-robotics task, used to simulate additional types of cognitive loading. This additional cognitive loading encourages participants to make use of the LoA switching functionality, and helps elucidate situations where the additional semantic information becomes useful. Participants are instructed to do their best on all tasks (equal importance). They conduct two trials comparing two conditions: semantics presented Fig. 1; semantics not presented Fig. 2. The sequence of two trials is counterbalanced among participants, to minimize the learning effects. Both trials share the same arena. However, victims’ locations are different, and starting/end points are swapped (clockwise/anti-clockwise path) between participants in every two trials.
Participants have five minutes for each trial. If they fail to bring the robot back to the endpoint within the 5-minute limit, the mission will be marked as a failure. Participants are not informed of the number of victims. They need to maximize their utilization of the 5 minutes. Hence, a time-to-completion metric is not meaningful. Instead, we examine success rates on each task.
The parallel task takes place continuously (endless task) during the mission. Participants need to identify whether a pair of 3D objects are identical, or are mirror images, see Fig. 3. This task has been validated in [25]. Standardized training [26] was provided on this task before the experiments to ensure all participants were aware of the effort required to manage this alongside the two robotics tasks.
Participants’ performances on this task (accuracy, completion, and completion speed), along with the robot-related tasks, are included in the overall evaluation. In addition, we recorded the baseline (accuracy of 10 questions) of each participant before the test and compared the baseline to the parallel task accuracy during the experiments. By comparing baseline performance with the performance during the robot-assisted missions, we can evaluate the impact of high-level semantics on participants’ parallel task execution quality.
@@ III-C Experimental design
A total of 16 participants (8 male / 8 female) were recruited from university staff and students aged 22-50, see LABEL:tab:_background. Most had a driver’s license or experience with 3D video games, which correlate somewhat with controlling remote vehicles [27]. Participants undertook standardised training to understand the system functionalities and control interface needed for the mission, including navigation in teleoperation and autonomy modes, robot’s SLAM map, and collision avoidance. Participants were introduced to high-level semantics, including SHA, radiation, and laser noise. When a semantic indicator score increases, the system emits a warning sound to attract human attention. Participants were trained how to react to situations, e.g. switch LoA to teleoperation when needed. Training is designed to avoid very large individual performance differences, especially if participants become lost in the arena.
TABLE I: Participants’ background before (first digit) and after data cleaning (digit after slash)
@@
Male 8/7 Driver only 5/5
@@
To minimize the effects of potential confounding factors, we surveyed participants’ background information LABEL:tab:_background and their feedback on arena and mission tasks. We removed 2 participants’ records from the database. One of them had a pre-existing bias toward the autonomy system, with no use of autonomy mode. She explained she had “zero trust in any robot system”. Another participant failed to follow the instructions and training given, which led to incomplete data. The second digit in LABEL:tab:_background refers to the number of participants after removal.
The UIs of the two experimental conditions are presented in Fig. 4 and Fig. 5.
Figure 4: UI for trials without semantic indicators: participants obtain SA from a pre-scanned lidar map, camera feed, real-time localisation, radiation readings, current robot LoA, and countdown of remaining mission time.
Figure 5: UI for trials with semantic indicators: three semantic indicator scores are displayed on the interface screen, in addition to the mapping and video data.
@@
The above training ensured that participants understood the system features and operation modes, creating a foundation for effective decision-making and navigation. Moreover, training gave participants an understanding of the workloads to help them make decisions how to utilise the variable autonomy and semantic indicator functionalities. To minimize the effects of various confounders, some additional information and tips were given to participants:
@@
  • The 5-minute time limitation is tight for most operators. Participants should carefully manage the time and areas to explore.
  • Try to follow the scheduled path (shortest navigation path) most of the time, and only deviate when participants feel it is necessary to achieve tasks.
@@
We used a Jackal mobile robot with a Velodyne 3D LiDAR, a Realsense D435i camera with an 87-degree horizontal view, and a Hamamatsu gamma ray sensor in the simulated arena. 3D LiDAR point clouds are transferred into a 2D LiDAR map. The robot navigated in the arena using customized AMCL and movebase ROS packages, with the same top speed limit in autonomy and teleoperation modes. Participants could switch LoA at any time.
@@ IV-A Significance analysis
Metric W Statistic P-value
Number of victims found 0.0 0.102
Parallel task accuracy 17.0 0.514
Parallel task accuracy degradation 17.0 0.515
Parallel task completion 42.0 0.542
Parallel task completion speed 31.0 0.194
Number of LoA switches 33.0 0.241
Teleoperation ratio 30.0 0.173
Trust in the system 10.5 0.080
Response time to switch LoA 16.0 0.040
NASA TLX 11.0 0.007
Situational trust scale 11.5 0.031
Navigation task completion 15.0 0.317
Given the number of participants and no assumption about the data, we conduct a significance analysis using the Wilcoxon signed-rank test, a non-parametric method. We calculate the W-statistic value and the P-value for each measurement. We consider a result statistically significant when the P-value is $<0.05$. We did not use Bonferroni or related p-value corrections as the study is exploratory.
We observed statistically significant differences in the reaction time to situation changes, the NASA TLX, and the situational trust scale. Table II shows that response time to switch LoA is significantly lower (W = 16.0, P = 0.040) with semantic indicators, i.e. high-level semantics significantly lower the reaction time to situation changes (e.g. a new detection of personal belongings). NASA TLX is significantly lower (W = 11.0, P = 0.007), meaning a lower perceived workload with the semantic indicators. The situational trust scale is significantly higher (W = 11.5, P = 0.031) with the semantics indicators, i.e. participants have more trust in their SA.

### v2

@@ Abstract
In this paper, we investigate the impact of high-level semantics (evaluation of the environment) on Human-Robot Team (HRT) and Human-Robot Interaction (HRI) in the context of mobile robot deployments. Although semantics has been widely researched in AI, how high-level semantics can benefit the HRT paradigm is underexplored, often fuzzy, and intractable. We applied a semantics-based framework that could reveal different indicators of the environment (i.e. how much semantic information exists) in a mock-up disaster response mission. In such missions, semantics are crucial as the HRT should handle complex situations and respond quickly with correct decisions, where humans might have a high workload. Especially when human operators need to shift their attention between robots and other tasks, they will struggle to build Situational Awareness (SA) quickly. The experiment suggests that the presented semantics: 1) alleviate the perceived human operator’s workload; 2) increase the operator’s trust in the SA; and 3) help to reduce the reaction time in switching the Level of Autonomy (LoA) when needed. Additionally, we find that participants with higher trust in the system are encouraged by high-level semantics to use teleoperation mode more.
Index Terms:
disaster response, situational awareness, semantics, cognitive workload, trust, level of autonomy
@@ I Introduction
The increasing use of robots is engendering challenges and higher demands for Human-robot collaboration. HRT needs to work together and leverage complementary strengths to achieve the shared mission goal by combining perception capabilities (remote sensing, object detection) and cognition (e.g. scene understanding, decision-making, task planning). SA [1] is one of the significant factors in bridging cognition and perception. Explainable and comprehensive SA can directly benefit robots and humans in many scopes, e.g. task planning [2], Variable Autonomy (VA) [3], and HRI.
@@
Semantic information, which is one of the components of SA, is suggested to alleviate the above difficulties [7]. Semantic information is widely discussed in the AI community, e.g. in computer vision. However, this literature is predominantly limited to relatively low-level semantics (e.g. image segmentation). The impact of high-level semantics, which requires more understanding and inference, has been less well-explored.
@@
Our previous work categorizes environmental semantics in the scope of disaster response [8] and proposes a framework [9] that applies the above categorization and quantifies different understandings of the situation by semantic indicators. Moreover, the framework enables providing robots with an indicator of the overall semantic richness of a scene. This method translates sensor readings and detections into a situational semantics richness score, which can inform human operators if the situation requires more attention.
@@
This paper provides new insights into HRT and HRI with respect to high-level semantics. We present three main contributions: i) identifying the usefulness of the proposed semantic richness framework [9]; ii) identifying how high-level semantics affect HRT and HRI in the context of a SAR mission; iii) insights about patterns and relationships among HRT factors and high-level semantics, e.g. LoA preference versus trust in a VA [10] system.
@@ II Related work
A wide range of studies explore ways of improving SA from the perspectives of visual design, workload and stress management, human training, and LoA.
Consequently, HRI can be categorized into cognitive responses, emotional responses, behavioral responses, and physiological responses [11]. Our work can partly be considered from a similar perspective, exploring how high-level semantics can enhance SA and benefit HRT and HRI, improving the likelihood of mission success, while reducing human workload.
@@
Figure 2: Test arena for trials without semantic indicators. Note that we need to use two different arena layouts for the two different trials, to minimise the confounder of learning effects between trials.
The above-mentioned studies investigated low-level semantics, but did not explore higher-level semantics. In contrast, our work explores high-level semantics, and its effects on human attention in HRT. [7] evaluated the effects of four interfaces on human stress and workload in a simulated fire surveillance and extinguishing mission. Virtual Reality interfaces were shown to enhance SA and reduce workload compared to conventional interfaces. The authors used high-level semantics, e.g. displaying information about fire risk, during the experiments. However, they investigated and compared the effects of different interface designs for displaying such data. In contrast, our work seeks to evaluate the impact of the high-level semantic data itself, i.e. performance with versus without displaying high-level semantics.
Other work explored how human training [15] and expertise affect SA. [16] investigated HRT under limited SA conditions, where operators lacked a bird’s-eye view of the environment. Participants were required to navigate robots until they met up together in the experiments. The authors emphasized the importance of training. Expert operators (several hours of training) outperformed others, by minimizing the number of commands issued to the robots. This highlights the critical role of prior knowledge and training in recovering and maintaining SA.
The LoA plays an important role in affecting SA as well. [17] explored human factors in HRT and found that high robot autonomy could reduce human SA. This makes it difficult for humans to regain SA when taking back control of the robot. The authors recommended a mixed-initiative approach that keeps humans in the loop and balances workload and SA between humans and robots. These related studies help us identify potential confounding factors and minimize their impacts in our experiment design.
@@
III Exploratory experiment of HRI and HRT with high-level semantics
This paper explores the effects of presenting high-level semantics in HRT deployments. Specifically, we explore the performance variance of HRT, with and without the display of semantic indicators (high-level semantics). We also explore the practical utility of our previously proposed framework for real-time computing of such indicators [9].
Our semantics framework [9] quantifies the richness of low-level semantics (salient features that can be understood by a human’s direct intuition) and high-level semantics (evaluation, or inference of a human’s perspective) [8] in the current scene. These high-level semantic indicators can be used to help a human operator achieve SA. In our experiments, high-level semantics include: SHA, radiation levels, and navigational laser noise level. SHA is detected as a distribution of personal belongings. Radiation level is a mapping from radiation sensor readings that enables human operators to visualise radiation intensity. The laser noise refers to the robot’s LiDAR. It helps robots and humans understand whether noise is becoming significant enough to disable the robot’s navigation. Each of these metrics carries a normalized score between 0 and 1, helping assess potential risks or challenges in real time.
In this exploratory study, we use a simulated SAR mission with human participants. Specifically, we simulate the first HRT deployment in a disaster response mission. The HRT needs to explore the environment and find victims while avoiding risks to robots. The robot can operate in two different LoAs: teleoperation and autonomous navigation modes. Humans can switch the LoA dynamically in a Human-Initiative paradigm [18] and can take over teleoperative control of the robot when they feel it is necessary. Moreover, parallel cognitive tasks are assigned to human operators to simulate non-robotic emergencies that may require human attention.
@@ III-A Arena
The experimental arena is an outdoor post-disaster environment, see Fig. 1 and Fig. 2. The robot is set to autonomously navigate along a pre-planned route, but the human operator can switch to teleoperation mode at any time, and explore other regions away from this route. Two human victims, one radiation source, and one laser noise zone are set up in the arena. Specifically, one human victim can be found by the camera (on the scheduled path, without the robot needing to deviate) by locating the personal belongings. Another victim is hidden behind obstacles, away from the scheduled path. The radiation source and laser noise zones are set to occur at pre-set regions along the planned path. The starting point and endpoint for robots are pre-defined. Note that we use a “within groups” experimental design, where each participant undertakes both experimental conditions (with and without semantic indicators). Therefore, we need to use different test arena layouts for each condition, to minimise learning effects as a confounder. However, the complexity of the tasks, and the types of semantics encountered, are kept very similar to ensure meaningful comparability.
@@ III-B Mission and Tasks
We assume that the HRT needs to conduct its first deployment to investigate the site in a SAR mission. The participant can switch back and forth between the teleoperation and autonomy. In autonomy mode, the robot will try to follow the pre-set route. Three tasks are assigned to the HRT: A) finding victims, B) navigating the robot, and C) parallel task. Task A requires the HRT to find as many victims as they can in the arena. Task B requires the HRT to navigate the robots from a pre-set starting point to the endpoint within set time limits. Task C is a non-robotics task, used to simulate additional types of cognitive loading. This additional cognitive loading encourages participants to make use of the LoA switching functionality, and helps elucidate situations where the additional semantic information becomes useful. Participants are instructed to do their best on all tasks (equal importance). They conduct two trials comparing two conditions: semantics presented Fig. 1; semantics not presented Fig. 2. To minimize the learning effects, the sequence of two trials is counterbalanced among participants. Moreover, victims’ locations are different in the two trials, and starting/end points are swapped (clockwise/anti-clockwise path) between participants.
Participants have five minutes for each trial. If they fail to bring the robot back to the endpoint within the time limit, the mission will be marked as a failure. Participants are not informed of the number of victims. They need to maximize their utilization of the time. Hence, a time-to-completion metric is not meaningful. Instead, we examine success rates on each task.
The parallel task takes place continuously (an endless task) during the mission. Participants need to identify whether a pair of 3D objects are identical, or are mirror images, see Fig. 3. This task has been validated in [19]. Standardized training [20] was provided on this task before the experiments to ensure all participants were aware of the effort required to manage this alongside the two robotics tasks.
Participants’ performances on this task (accuracy, completion, and completion speed), along with the robot-related tasks, are included in the overall evaluation. In addition, we recorded the baseline (accuracy of ten questions) of each participant before the experiment and compared the baseline to the parallel task accuracy during the experiments. From this comparison, we can evaluate the impact of high-level semantics on participants’ parallel task execution quality.
@@ III-C Experimental design
A total of 16 participants were recruited from university staff and students aged 22-50, see Table I. Most had a driver’s license or experience with 3D video games, which correlate with controlling remote vehicles [21]. Participants undertook standardised training to understand the system functionalities and control interface needed for the mission, including navigation in teleoperation and autonomy modes, robot’s SLAM map, and collision avoidance. Participants were introduced to high-level semantics, including SHA, radiation, and laser noise. When a semantic indicator score increases, the system emits a warning sound to attract human attention. Participants were trained on how to react to situations, e.g. switch LoA to teleoperation when needed. Training is designed to avoid very large individual performance differences, especially if participants become lost in the arena.
TABLE I: Participants’ gender and background
@@
Male 8/7* Driver only 5/5
@@
  • * Red digits refer to the number of participants after data cleaning, which are involved in the future analysis.
To minimize the effects of potential confounding factors, we surveyed participants’ background information in Table I and their feedback on arena and mission tasks. We removed two participants’ records from the database. One had a pre-existing bias toward the autonomy system, with no use of autonomy mode. She explained she had “zero trust in any robot system”. Another one failed to follow the instructions given, which led to incomplete data. The red digit in Table I refers to the number of participants after removal.
The GUIs of the two experimental conditions are presented in Fig. 4 and Fig. 5.
Figure 4: GUI for trials without semantic indicators: participants obtain SA from a pre-scanned LiDAR map, camera feed, real-time localisation, radiation readings, current robot LoA, and countdown of remaining mission time. The bubble is used for clearly presenting the semantic indicator scores. It does not show up in the GUI during the experiments.
Figure 5: GUI for trials with semantic indicators: three semantic indicator scores are displayed on the interface screen, in addition to the mapping and video data.
@@
The above training ensured that participants understood the system features and operation modes, creating a foundation for effective decision-making and navigation. Moreover, training gave participants an understanding of the workloads to help them make decisions on how to utilise the VA and semantic indicator functionalities. To minimize the effects of various confounders, some additional information and tips were given to participants:
@@
  • The time limitation is tight for most operators. Participants should carefully manage the time and areas to explore. For the same reason, try to follow the scheduled (shortest) path most of the time, and only deviate when participants feel it is necessary.
@@
We used a Jackal mobile robot with a Velodyne 3D LiDAR, a Realsense D435i camera, and a Hamamatsu gamma ray sensor in the simulated arena. The robot navigated in the arena using customized AMCL and movebase ROS packages, with the same top speed limit in autonomy and teleoperation modes. Participants could switch LoA at will.
@@ IV-A Significance analysis
Metric $W$ $P$ $r$
Number of victims found 0.0 0.102 0.436
Parallel task accuracy 17.0 0.514 0.174
Parallel task accuracy degradation 17.0 0.515 0.174
Parallel task completion 42.0 0.542 0.163
Parallel task completion speed 31.0 0.194 0.347
Number of LoA switches 33.0 0.241 0.313
Teleoperation ratio 30.0 0.173 0.365
Trust in the system 10.5 0.080 0.468
Response time to switch LoA 16.0 0.040 0.570
NASA TLX 11.0 0.007 0.724
Situational trust scale 11.5 0.031 0.577
Navigation task completion 15.0 0.317 0.267
Given the number of participants and no assumption about the data, we conduct a significance analysis using the Wilcoxon signed-rank test, a non-parametric method. We calculate the W-statistic value $W$, the P-value $P$, and the effect size $r$ for each measurement. We consider a result statistically significant when $P$ is <0.05<0.05, and a meaningful and strong practical difference when $r$ is >0.5>0.5 according to Cohen’s standard [23]. We did not use Bonferroni or related p-value corrections as the study is exploratory.
We observed statistically significant differences in the reaction time to situation changes, the NASA TLX, and the situational trust scale. Table II shows that response time to switch LoA is significantly lower and shows a strong practical difference ($W$ = 16.0, $P$ = 0.040, $r$ = 0.570) with semantic indicators. It means high-level semantics significantly help to lower the reaction time to situation changes (e.g. a new detection of personal belongings). NASA TLX is significantly lower and shows a strong practical difference ($W$ = 11.0, $P$ = 0.007, $r$ = 0.724). It refers that semantic indicators lower the perceived human workload. Moreover, the situational trust scale is significantly higher and shows a strong practical difference ($W$ = 11.5, $P$ = 0.031, $r$ = 0.577) with the semantic indicators. It means participants have more trust in their SA with the assistance of semantic indicators.